OpenStudy (anonymous):

solve the definite integral from 0 to pi/2 of cos^2(x)dx

OpenStudy (tkhunny):

What's preventing you from solving it?

OpenStudy (anonymous):

i'm suppposed to evaluate it using the integral from 0to pi/2 of cosx dx being equal to the integral from 0 to pi/2 of sinx dx. I dont see how this relates to cos^2(x)dx

OpenStudy (tkhunny):

That's odd. It's a lot easier if you let some trigonometry help you. Think in this important identity: \(\cos(2x) = 2\cdot\cos^{2}(x) - 1\) Are we insisting that we use the given method?

OpenStudy (anonymous):

yeah because its a proof :(

OpenStudy (tkhunny):

You can integrate by parts, once, and prove the arguments \(\cos^{2}(x)\;and\;\sin^{2}(x)\) produce identical results.

OpenStudy (anonymous):

ok, im gonna give it a try

OpenStudy (tkhunny):

Split it up cos(x) and cos(x).

OpenStudy (anonymous):

yeah I think i got it. thank you

OpenStudy (tkhunny):

Super! Good work.

OpenStudy (anonymous):

i have to evaluate and i just evaluated them both to equal 1/2

OpenStudy (tkhunny):

?? It should be "1" with just cosine and pi/4 with cosine squared.

OpenStudy (tkhunny):

I'm still not sure how we used the hint.